please show explanation for this problems
1)if f(x) = ax+b, f(0) = 2, and f(1)=-2
2} The set {x:x is an integer and 1x - 11 < 2} equals ?
3) The inverse of the function y = 2x is ?

Answer by rothauserc(4718)

****************************
1) f(0)=2 implies b=2
f(1)=-2 then a+2=-2 and a=-4
therefore f(x) = -4x+2
****************************
2) all x < 13
****************************
3) y = 2x
x = y/2
y = x/2
f^(-1) = x/2
****************************